# Theory of Code Space: Do Code Agents Understand Software Architecture?

### 28 Feb 2026 v1

@@ 7.1 Implications for Code Agent Design
Current code agents (SWE-agent, Aider, Claude Code, Cursor) use tool-augmented retrieval to navigate repositories, but none explicitly construct or maintain a structured architectural belief. Our results show that the strongest LLM agents can surpass rule-based strategies when they combine semantic comprehension with reliable belief externalization (Claude Sonnet 4.6), but that most models fail at the externalization step despite demonstrating code understanding in their exploration behavior.
This suggests three improvement paths: (1) hybrid approaches combining AST-level import extraction with LLM semantic analysis, ensuring structural completeness while adding semantic depth; (2) belief externalization training—explicitly optimizing models to faithfully serialize architectural knowledge into structured formats; and (3) exploration strategy optimization—models like Gemini 3.1 Pro that under-explore (opening only 9 files) could improve significantly with better budget allocation.
@@ 7.2 The Belief Externalization Problem
The spread across models is striking: Claude Sonnet 4.6 recalled 55% of IMPORTS edges with 100% precision, while Gemini 3 Flash recalled only 9% despite opening a similar number of files. This 6 gap in IMPORTS recall—for edges the models demonstrably saw—suggests that belief externalization capability varies dramatically across model families, not just between LLMs and baselines.
Interestingly, Claude Sonnet 4.6’s near-perfect precision (0.991) across all edge types shows that faithful externalization is achievable with current models. The challenge appears model-specific rather than fundamental: it is not that structured JSON is inherently difficult to produce, but that some models have not been optimized for this capability. Future work should investigate whether fine-tuning, chain-of-thought prompting, or iterative self-checking can close this gap for weaker models.
@@ 7.3 Limitations
ToCS v0.1 is deliberately narrow in scope. Generated codebases, while structurally realistic, may not capture all patterns of real-world architecture. In particular, our use of neutral filenames (mod_a.py rather than authentication_service.py) removes naming priors that real codebases provide, making the task harder in one dimension while keeping it easier in others (clean structure, no organic complexity, limited dead code). This design choice isolates architectural reasoning from lexical guessing, but means our difficulty estimates may not transfer directly to real repositories. We evaluate a single architectural pattern (Pipeline) in a single language (Python). Our LLM results are preliminary: five models from three providers, each evaluated on 3 codebases with a single run per model-codebase pair (no variance estimation across runs). GPT-5.3-Codex’s high cross-codebase variance (F1 ranging from 0.392 to 0.719) highlights the need for repeated runs. The cognitive map externalization is a proxy for internal representation—a model may “know” more than it serializes, as the wide variation across models suggests. We plan to validate via correlation between map quality and counterfactual probe accuracy, and to expand to additional models (open-weight), architectural patterns, repeated runs with variance estimation, and the REVISE evaluation mode.
7.4 Community Call
We release the complete ToCS toolkit—codebase generator, partial observability harness, evaluation pipeline, and scoring metrics—as open-source software. We invite the community to:
1.
Evaluate additional models: Run open-weight models (Llama, Qwen, DeepSeek) and replicate our frontier model results with repeated runs.
2.
Contribute codebase patterns: Extend beyond Pipeline to event-driven, microservices, plugin systems, hexagonal architecture (see CONTRIBUTING.md).
3.
Add languages: The harness is language-agnostic; the generator needs new templates for TypeScript, Go, etc.
4.
Test scaffold-augmented agents: Evaluate whether static analysis tools, RepoMap-style context, or explicit state tracking improve map quality.
@@ 8 Conclusion
We introduced ToCS, a benchmark for evaluating whether code agents can construct coherent architectural beliefs through active codebase exploration under partial observability. Our procedural generator produces medium-complexity codebases with four typed edge categories and planted invariants. Preliminary experiments with four baselines and five frontier LLM agents from three providers reveal that: (1) exploration strategy limits discovery for rule-based methods, but the strongest LLM agent (Claude Sonnet 4.6, F1 = 0.646) surpasses all baselines by combining semantic comprehension with reliable belief externalization; (2) LLM agents collectively discover all four edge types—including DATA_FLOWS_TO edges (GPT-5.3-Codex, 31% recall) and REGISTRY_WIRES edges (Claude Sonnet 4.6, 100% recall)—capabilities entirely absent from rule-based methods; yet (3) weaker LLM agents score below simple heuristics due to severe belief under-reporting, with a 6 spread in IMPORTS recall across models (9–57%), confirming that structured belief externalization remains a key capability gap. The 35-point F1 gap between Oracle and the best LLM agent, combined with zero invariant discovery across all methods, defines a challenging and discriminative evaluation space. We release ToCS as an open benchmark and invite the community to replicate and extend these preliminary results.
@@ Appendix C Budget Sweep
At tight budgets (), Config-Aware’s strategy of prioritizing registry files yields a 3 advantage over Random (0.175 vs. 0.056). As budget increases, raw coverage dominates: at (allowing agents to open nearly all files), Random matches Config-Aware (0.632 vs. 0.626). BFS-Import shows the steepest growth from to (0.293 0.603), reflecting its systematic but slow import-chain traversal. The convergence at high budgets confirms that the current medium-complexity codebases require tighter budgets or larger codebases to strongly discriminate exploration strategy for structural edge types; the meaningful discrimination at lies in semantic edge discovery by LLM agents.

### 05 Mar 2026 v3

@@ 7.1 Implications for Code Agent Design
Current code agents (SWE-agent, Aider, Claude Code, Cursor) use tool-augmented retrieval to navigate repositories, but none explicitly construct or maintain a structured architectural belief. Our results show that the strongest LLM agents can surpass rule-based strategies when they combine semantic comprehension with reliable belief externalization (GPT-5.3-Codex and Claude Sonnet 4.6 both exceed all baselines), but that weaker models fail at the externalization step despite demonstrating code understanding in their exploration behavior.
This suggests four improvement paths: (1) hybrid approaches combining AST-level import extraction with LLM semantic analysis, ensuring structural completeness while adding semantic depth; (2) belief externalization training—explicitly optimizing models to faithfully serialize architectural knowledge into structured formats; (3) exploration strategy optimization—our APG decomposition (§6.4) shows a 6–17-point gap between oracle and agent file selection across models, confirming that which files to open is a key bottleneck; and (4) explicit state management—our probe ablation (§6.5) shows that retaining structured belief maps in context yields a 14-point F1 boost for GPT, empirically validating the value of persistent, accumulative state structures—though the effect is model-dependent, indicating that self-scaffolding is itself a capability that varies across models.
@@ 7.2 The Belief Externalization Problem
GPT-5.3-Codex recalled 69% of IMPORTS edges while Gemini 3 Flash and 2.5 Pro recalled only 11–12% despite opening similar numbers of files. This 6 gap—for edges the models demonstrably saw—shows that belief externalization capability varies dramatically across model families. Both GPT and Claude demonstrate that faithful externalization is achievable: the challenge is model-specific, not fundamental.
Belief state instability.
The Gemini family reveals three distinct instability patterns. Gemini 2.5 Pro shows catastrophic collapse: building a map by step 9 (peak F1 = 0.33), then losing 12 correct edges in a single probe. Gemini 3 Flash shows recency bias: each probe reports only 3–5 recently-examined components. Gemini 3.1 Pro shows variable instability: monotonic accumulation on its best codebase but collapse to zero edges on its worst. Critically, Gemini 2.5 Flash—the smallest model—exhibits zero correct-edge loss across all probes. This suggests belief state maintenance depends on training objectives, not model scale, and may stem from treating each probe as “summarize from scratch” rather than “incrementally update.” Externalization quality is also likely prompt-dependent; future work should include prompt ablation studies to disentangle model capability from prompt sensitivity.
Error analysis: prompt ambiguity drives false positives.
Of Gemini 2.5 Flash’s 40 false positive edges, only 2 (5%) are true hallucinations. The dominant failure (45%) is edge type confusion: reporting both IMPORTS and CALLS_API when code both imports and calls a symbol—arguably reasonable given underspecified definitions. Another 22% come from treating non-component files as edge endpoints. This reveals that prompt specification is a first-order variable: what looks like an externalization failure may be a legitimate alternative interpretation.
| Edge | Gemini 2.5 Flash | Claude Sonnet 4.6 |
| cli config | IMPORTS | IMPORTS |
| cli runner | IMPORTS | IMPORTS |
| cli config::PipelineConfig | CALLS_API (FP) | — |
| cli runner::run_pipeline | CALLS_API (FP) | — |
| cli runner | — | CALLS_API (TP) |
7.3 Benchmark Design Lessons
Building ToCS v0.1 exposed several pitfalls in belief-probing benchmark design:
L1: Edge type decision rules. When code both imports and calls a symbol, the prompt must specify which edge type to report. Ambiguity here caused 45% of Gemini 2.5 Flash’s false positives. Fix: explicit decision trees.
L2: Component boundaries. Without defining whether test files or configs count as “components,” 22% of FPs came from including non-component endpoints. Fix: explicit inclusion rules in the probe.
L3: Invariant field semantics. Generic fields without per-type definitions yielded Inv F1 = 0.0 for all models. Adding worked examples jumped scores to 0.52–0.78 under relaxed matching.
L4: Alternative correct answers. Strict exact-match scoring penalizes semantically correct but differently-formatted answers. Fix: tiered or normalized scoring.
L5: Prompt specification as experimental variable. What looks like a model capability gap may be a prompt specification gap. A single prompt change moved Gemini 2.5 Flash’s F1 by 0.141 while barely affecting Claude (0.012; Table 8 in Appendix).
L6: Probes as active interventions. In scratchpad mode, improved probe prompts change not just serialization quality but exploration behavior—on one codebase, GPT-5.3-Codex opened 4 additional stage modules after a more structured probe revealed gaps in its own map (F1: 0.392 0.636). The probe condition ablation (§6.5) quantifies this: scratchpad boosts GPT by 14 F1 points but has no dependency-level effect on Gemini—suggesting different models leverage self-scaffolding through different channels.
@@ 7.4 Limitations
ToCS v0.1 evaluates a single architectural pattern (Pipeline) in a single language (Python) with neutral filenames (mod_a.py), no standalone documentation, and no organic complexity. Six models from three providers are each evaluated on 3 codebases with a single run per model-codebase pair. All results use a single prompt design; different formulations could yield substantially different rankings. The cognitive map externalization is a proxy for internal representation—a model may “know” more than it serializes.
7.5 Roadmap
Phase 1 (current): benchmark construction—stress-test the framework with diverse models. Phase 2: community-contributed codebase patterns, additional languages, documentation as a discovery channel, and REVISE mode evaluation. Phase 3: large-scale controlled experiments with repeated runs, open-weight models, and prompt ablation studies. We release the complete toolkit as open-source and invite contributions.
@@ 8 Conclusion
We introduced ToCS, the first benchmark for active architectural belief construction in code. Three findings stand out. First, the Active-Passive Gap is model-dependent: GPT-5.3-Codex builds better maps through active exploration () while Gemini 2.5 Flash does better receiving all files at once ()—showing that active exploration is itself a non-trivial capability. Second, self-scaffolding through belief externalization is model-dependent: scratchpad mode boosts GPT by 14 F1 points but not Gemini. Third, belief state maintenance varies dramatically: the smallest Gemini model maintains perfectly stable beliefs while its larger sibling suffers catastrophic collapse. These findings validate the benchmark’s design—periodic probing reveals dynamics invisible to final-snapshot evaluation—while the design lessons learned from building it (§7.3) are themselves a contribution. We release ToCS as an open benchmark and invite community contributions.
@@ Appendix C Budget Sweep
At tight budgets (), Config-Aware’s strategy of prioritizing registry files yields a 3 advantage over Random (0.175 vs. 0.056). As budget increases, raw coverage dominates: at , Random matches Config-Aware. The meaningful discrimination at lies in semantic edge discovery by LLM agents.
Appendix D Prompt Sensitivity
| Model | Old Prompt | New Prompt | F1 |
| Gemini 2.5 Flash | 0.328 | 0.469 | +0.141 |
| GPT-5.3-Codex | 0.564 | 0.676 | +0.113 |
| Claude Sonnet 4.6 | 0.639 | 0.650 | +0.012 |
Models with weaker baseline externalization benefit disproportionately from prompt improvements. Claude Sonnet 4.6, which already externalized comprehensively under the old prompt, showed minimal change (0.012). Main-text results use the improved prompt throughout.